Larkspur Furniture Company started construction of a combination office and warehouse building for its own use at an estimated cost of $5,004,000 on January 1, 2017. Larkspur expected to complete the building by December 31, 2017. Larkspur has the following debt obligations outstanding during the construction period. Construction loan-12% interest, payable semiannually, issued December 31, 2016 $1,990,900 Short-term loan-10% interest, payable monthly, and principal payable at maturity on May 30, 2018 1,599,600 Long-term loan-11% interest, payable on January 1 of each year. Principal payable on January 1, 2021 995,500Assume that Larkspur completed the office and warehouse building on December 31, 2017, as planned at a total cost of $5,151,300, and the weighted-average amount of accumulated expenditures was $3,764,600. Compute the avoidable interest on this project. (Use interest rates rounded to 2 decimal places, e.g. 7.58% for computational purposes and round final answers to 0 decimal places, e.g. 5,275.) Avoidable Interest $Compute the depreciation expense for the year ended December 31, 2018. Larkspur elected to depreciate the building on a straight-line basis and determined that the asset has a useful life of 30 years and a salvage value of $297,500. (Round answer to 0 decimal places, e.g. 5,275.) Depreciation Expense $
Answer #1

Solution:

1.)

Avoidable interest:

Weighted average

Accumulated Expenditure × Interest Rate = Avoidable Interest
$1,990,900 × 12% = $238,908
$1,386,700 ($5,151,300 - $3,764,600) × 10.38% = $143,939
$3,377,600 $382,847

Weighted average interest rate computation

Principal Interest
10% short term loan $1,599,600 $159,960 ($1,599,600 ×10%)
11% long term loan $995,500 $109,505 ($995,500 × 11%)
Total $2,595,100 $269,465
Interest rate Total interest / Total principal $269,465 / $2,595,100 = 10.38%

2.)

Actual interest

Construction loan $1,990,900 × 12% = $238,908
Short term loan $1,599,600 × 11% = $175,956
Long term loan $995,500 × 10% = 99,550
Total $514,414

We use avoidable interest, because avoidable interest is lower than actual interest.

Cost $5,151,300
Interest captialized $382,847
Total cost $5,534,147

Depreciation expenses =$5,534,147 -$297,500 / 30 years

=$5,236,647/30 years

=$174,554.90 or $174,555